OpenStudy (anonymous):

sq root of 20x times sq rt. 30x^3 is 6x sq rt of 5 yes or no?

OpenStudy (anonymous):

sq rt. 30x^3= 30x^3/2

OpenStudy (anonymous):

square root (20 * 30x^3/2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

square root (20 * 30x^3/2)= 20 ^1/2 * 30x^3/4

OpenStudy (apoorvk):

what do you think? yes or no?

OpenStudy (lgbasallote):

\[\huge \sqrt{20x} * \sqrt {30x^{3}} = 6x\sqrt{5}\] is the question right? so... \[\huge \sqrt{(20x)(30x^{3})} = 6x \sqrt{5}\] \[\huge \sqrt{(600x^{4})} = 6x\sqrt{5}\] \[\Large \sqrt{(2)(3)(10)(10)(x)(x)(x)(x)} = 6x \sqrt{5}\] factor out of the square root all that appeared twice... \[\large (10)(x)(x)\sqrt{(2)(3)} = 6x\sqrt{5}\] \[10x^{2}\sqrt{6} \neq 6x\sqrt{5}\]